OpenStudy (anonymous):

It's a very simple question, and it's not even for homework. Well, technically it is for homework, but it's not the homework question. I just forgot what this was. Question: What is the difference of squares formula? Again, not a homework question, I literally just forgot what it was. Thanks for helping!

OpenStudy (anonymous):

@sleepyjess @ganeshie8 @mathmale @mathstudent55 @zepdrix @Mehek14

OpenStudy (anonymous):

HOORAY! @sleepyjess IS HERE!

OpenStudy (sleepyjess):

Ah, let me find this in my notes

OpenStudy (anonymous):

K, thx :)

OpenStudy (sleepyjess):

\(a^2 - b^2 = (a – b)(a + b)\)

OpenStudy (anonymous):

Alright. So for (x^2 - 9), the difference of squares would be: (x^2 - 3^2) = (x - 3)(x + 3), correct @sleepyjess ?

OpenStudy (sleepyjess):

And as proof, \(4^2 - 2^2 = \\(4-2)(4+2)\\(2)(6)\\12\\4^2 - 2^2 =\\16 - 4\\12\)

OpenStudy (sleepyjess):

Yep :)

OpenStudy (anonymous):

HOORAY! Thanks bro :) @sleepyjess IS AWESOME!

OpenStudy (sleepyjess):

haha no problem :) happy to help

OpenStudy (anonymous):

:)